OpenStudy (anonymous):

A phone company offers two monthly plans. Plan A costs $29 plus an additional $.11for each minute of calls. Plan B costs $11plus an additional $0.15 for each minute of calls. For what amount of calling do the two plans cost the same? What is the cost when the two plans cost the same?

OpenStudy (texaschic101):

29 + 0.11m = 11 + 0.15m --- subtract 0.15m from both sides 29 + 0.11m - 0.15m = 11 --- subtract 29 from each side 0.11m - 0.15m = 11 - 29 --- combine like terms -0.04m = - 18 m = -18/-0.04 m = 450 They are equal at 450 minutes 29 + 0.11m = 29 + 0.11(450) = 29 + 49.50 = 78.50 11 + 0.15m = 11 + 0.15(450) = 11 + 67.50 = 78.50 and at 450 minutes, each plan costs $ 78.50